OpenStudy (anonymous):

Which speed is the fastest (use 1 mile ≈ 1.6 kilometers)?Select one of the options below as your answer:A. 12 miles/hourB. 18 feet/secondC. 19 kilometers/hourD. cannot be determined

OpenStudy (ghazi):

12miles= 12*1.65 kms/hr , 18 feet= 18/3.2 meters per sec= 18/3.2*18/5 kms/hr, and 19 kms per hour..now you can check ..which is the fastest one

OpenStudy (anonymous):

No, I don't under stand a word you just said

OpenStudy (anonymous):

is it 9219 km /hr?

OpenStudy (ghazi):

i just stated a bit of conversion in kms/hour and after solving you will get 18 feet/s as the fastest one since it is equal to 20.5kms/hr

OpenStudy (anonymous):

is it 19 km/ hr?

OpenStudy (ghazi):

no it's 18 feet/ sec